Question 2 Tries remaining: 2 Points out of 7.70 P Flag question Suppose the demand for jackets was given by: Qp- 190 -0.2P. The supply of jackets is given by: Qs- 2P -30. Suppose the price was $34 per jacket. Calculate whether there is a surplus or shortage of jackets at that price and the quantity of jackets associated with the surplus or shortage. Enter your response as Qs Negative numbers indicate shortage. Positive numbers indicate surplus. (Round to the nearest two decimal places if necessary.) Answer: Check
Answer #1

If P= 34,

Qd = 190-0.2P = 190-0.2(34) = 183.2

Qs = 2P-30 = 2(34)-30 = 68-30 = 38

Since Qd>Qs, so there is a shortage of 183.2 - 38 = 145.20

So, your answer will be - 145.20. (Don't forget to put minus along with your answer)